provide Safe Doses calculation, why patient receive this? nursing implication/interaction/teaching. lease type :)
Answer #1

There are 3 primary methods for calculation of medication dosages; Dimensional Analysis, Ratio Proportion, and Formula or Desired Over Have Method. We are going to explore the Desired Over Have or Formula Method, one of these 3 methods, in more detail.

Desired Over Have or Formula Method uses a formula or equation to solve for an unknown quantity (x) much like ratio proportion.

Drug calculations require the use of conversion factors, for example when converting from pounds to kilograms or liters to milliliters. Simplistic in design, this method affords clinicians the opportunity to work with various units of measurement, converting factors to find the answer. These methods are useful in checking the accuracy of the other methods of calculation, thus acting as a double or triple check.

Technique

There are 3 primary methods for the calculation of medication dosages as referenced above. These include Desired Over Have Method or Formula, Dimensional Analysis and Ratio and Proportion (as cited in Boyer, 2002)[Lindow, 2004].

Desired Over Have or Formula Method

Desired over Have or Formula Method is a formula or equation to solve for an unknown quantity (x) much like ratio proportion. Drug calculations require the use of conversion factors, such as when converting from pounds to kilograms or liters to milliliters. Simplistic in design, this method affords us the opportunity to work with various units of measurement, converting factors to find our answer. Useful in checking the accuracy of the other methods of calculation as above mentioned, thus acting as a double or triple check.

  • A basic formula, solving for x, guides us in the setting up of an equation:
  • D/H x Q = x, or Desired dose (amount) = ordered Dose amount/amount on Hand x Quantity.

For example, a provider requests lorazepam 4 Mg IV Push for a patient in severe alcohol withdrawal. The clinician has 2 mg/mL vials on hand. How many milliliters should he or she draw up in a syringe to deliver the desired dose?

  • Dose ordered (4 mg) x Quantity (1 mL)/Have (2 mg) = Amount wanted to give (2 mL)

Units of measurement must match, for example, milliliters and milliliters, or one needs to convert to like units of measurement. In the example, above, the ordered dose was in milligrams, and the have dose was in milligrams, both which cancel out leaving milliliters (answer called for milliliters), so no further conversion is required.

Dimensional Analysis Method

An order placed by a provider for lorazepam 4 mg IV PUSH for CIWA score of 25 or higher, follow CAGE Protocol for subsequent dosages based on CIWA scoring.

  • The clinician has 2 mg/mL vials in the automated dispensing unit.
  • How many milliliters are needed to arrive at ordered dose?
  • The desired dose os placed over 1 remember, (x mL) = 4 mg/1 x 1 mL/2 mg x (4)(1)/2 x 4/2 x 2/1 = 2 mL, keep multiplying/dividing until the desired amount is reached, 2 mL in this example.
  • Notice, the fraction was set up with milligrams and milligrams strategically placed so like units could cancel each other out, making the equation easier to solve for the unit desired or milliliters. The answer makes sense, so work is done.

Zeros can be canceled out in the same way as like units. For example:

  • 1000/500 x 10/5 = 2, the 2 zeros in 1000 and 2 zeros in 500 can be crossed out since like units in numerator and denominator, leaving 10/5, a much easier fraction to solve and the answer makes sense.

We have addressed zeros, and now let us look at 1.

  • If one multiplies a number by a 1, then the number is unchanged.
  • In contrast, if you multiply a number by zero, the number becomes zero.
  • Examples listed below are as follows: 18 x 0 = 0 or 20 x 1 = 20.

Ratio and Proportion Method

The Ratio and Proportion Method has been around for years and is one of the oldest methods utilized in drug calculations (as cited in Boyer, 2002)[Lindow, 2004]. Addition principals is a problem-solving technique that has no bearing on this relationship, only multiplication, and division are used to navigate through a ratio and proportion problem, not adding. An example listed below will provide a better explanation using a fraction or a colon format:

A provider orders lorazepam 4 mg IV Push now for a CIWA score of 25. There are 2 mg/mL vials on hand. How many milliliters are required to carry out the ordered dose?

  • Have on hand / Quantity you have = Desired Amount / x
  • 2 mg/1 mL = 4 mg/x
  • 2x/2 = 4/2
  • x = 2 mL

In colon format, one would use H:V::D:X and multiply means DV and Extremes HX.

  • Hx = DV, x = DV/H, 2:1::4:x, 2x = (4)(1), x = 4/2, x = 2 mL

2016 study evaluated the role confidence plays in overall arithmetic in drug calculation skills. Study participants attended remedial math classes from a wide range of educational backgrounds and age dynamics seeking a first degree in nursing, a foundation degree, or post-registration courses (Shelton, 2016). The study revealed one-third of students feel a lack of confidence which originated in an earlier stage of education dating back to a primary school environment (Shelton, 2016). The study concluded that confidence plays a role in dosage calculations and overall performance of mathematical calculations and can be improved in an environment that fosters a deep-learning approach (Shelton, 2016).